OpenStudy (anonymous):

Write an equation of a line whose y-intercept is 2 and whose slope is -1

OpenStudy (anonymous):

y=mx + b is the standard equation for a line with slope m and y intercept b. so it's just a matter of plugging in now. y = -1x + 2 y = -x + 2

OpenStudy (anonymous):

fyeah! thanks!